# Learning a Neural Solver for Parametric PDE to Enhance Physics-Informed Methods

### v2

@@ B.3 Optimality and Convergence
B.4 Conclusion
As a conclusion of this proof sketch, we showed that, under optimality conditions and with linearization of the neural network, we have:
  1. (i)
@@ B.4 Conclusion
This analysis underscores the effectiveness of using neural networks within the Physics-Informed framework, leveraging their ability to reduce the condition number and improve convergence characteristics.
Appendix C Implementation details
C.1 B-Spline basis
Figure 3: B-spline basis with $N=10$ terms with shifted spline (Left) and higher multiplicity nodes (Right). Dashed lines represent nodes position with color the darker, the higher the multiplicity.
Higer-dimension basis
For $1$d + time, $2$d dataset and $2$d + time we build a $2$ (or $3$) dimensional B-spline basis, i.e., we treat the time coordinate as a variable as spatial ones. To build such bases, we compute the Cartesian product between the $2$ (or $3$) $1$d-bases, $1$ per dimension. This means that for a $2$d dataset, for which we chose to use bases of size $N1$ and $N2$ for each coordinate, the resulting basis will have $N1+N2+N1*N2$ terms. This method makes the training more costly and several techniques to improve its scalability could be used. For $3$d datasets, the number of terms in the basis is cubic.
C.2 Training details
In our experiments, neural networks are trained using the Adam optimizer. For network optimization, we employ a smooth $l1$-loss for our solver while for other baselines, we use MSE loss and/or physical losses. All models are trained for at least $1,500$ epochs on datasets composed of some sampling of $\gamma$ and/or $f$ and/or $g$. If not stated otherwise, we train our model for $750$ epochs and baselines for $1,500$ epochs, with Adam and an initial learning rate of $0.001$. We use an exponential learning rate scheduler that lowers the learning rate by $0.995$ every epoch. Experiments are conducted on NVIDIA TITAN V ($12$ Go) for $1$d datasets and NVIDIA RTX A6000 GPU with 49Go for $1d$ + time, $2$d datasets or $2$d + time. We show on algorithm 3 a pseudo-code for training of our proposed method.
Data: $\theta_{0}\in\mathbb{R}^{n}$, PDE ($\gamma,f,g)$
for $e=1...$ epochs do
       for (PDE, x, u) in dataset do
             Initialize $\theta_{0}$
             Compute $\theta^{L}=\textnormal{Solve}(\theta_{0},\gamma,f,g)$ using Algorithm 1
             Reconstruct $u_{\theta^{L}}(x)$
             Update the parameters $\nu$ of $\mathcal{F}$ with gradient descent from the data loss in Equation 10.
       end for
end for
C.3 Models
We present here the architecture details for our model and the baseline on each dataset. In all our experiments, we use a GeLU (Hendrycks & Gimpel, 2016) activation function. The details of the model architecture on each dataset are presented in table 3.
Model Architecture Dataset
1d 1d + time 2d 2d + time
Helmholtz Poisson NLRD Darcy Heat
PINNs MLP depth 3 3 3 3 3
MLP width 256 256 256 256 256
learning rate 1e-4 1e-5 1e-4 1e-3 1e-3
training steps 1000 1000 1500 1500 1500
PPINNs MLP depth 8 3 5 3 3
MLP width 64 256 256 256 256
PO-DeepONet Branch Net depth 3 2 5 5 5
Branch Net width 256 256 256 256 256
Trunk Net depth 3 2 5 5 5
Trunk Net width 256 256 256 256 256
MLP + basis MLP depth 5 5 5 5 5
MLP width 256 256 1,024 1,024 1,024
PI-DeepONet Branch Net depth 3 2 5 5 5
Branch Net width 256 256 256 256 5
Trunk Net depth 3 2 5 5 5
Trunk Net width 256 256 256 256 256
PINO FNO depth 3 3 4 3 3
FNO width 64 64 64 64 64
FNO modes 1 16 16 10 20 7
FNO modes 2 - - 5 20 7
FNO modes 3 - - - - 5
FNO fc dim 64 64 64 64 64
Ours FNO depth 3 3 3 3 3
FNO width 64 64 64 64 64
FNO modes 1 16 16 10 20 7
FNO modes 2 - - 5 20 7
FNO modes 3 - - - - 5
FNO fc dim 64 64 64 64 64
Table 3: Architecture details of our model and baselines
Hyperparameters for Helmholtz:
Training of PPINNs:
For the parametric PINN baseline, we considered different training schemes as using a unique optimizer or an improved training scheme from (Krishnapriyan et al., 2021), i.e. using mixed optimization with Adam and L-BFGS to train the model. For the Helmholtz dataset, we trained our model for 2000 epochs with a plateau learning rate scheduler with patience of 400 epochs. For the Poisson dataset, we trained our parametric PINN model for 5000 epochs with a cosine annealing scheduler with a maximum number of iterations of 1000. For the Reaction-diffusion dataset, we consider an initial learning rate of $0.0001$ instead of $0.001$.
PINNs baseline for Preconditioning:
The conditioning of the problem highly depends on the parameters of the PDE and initial/boundary conditions. This can lead to unstable training when optimizing PINNs, requiring specific parameters configuration for each PDEs. To avoid an extensive research of the best training strategy, we found a configuration that allowed a good fitting of most of the testing dataset. This means that in the values reported in table 2, we removed trainings for which the losses exploded (only a few hard PDEs were removed, typically between none to $20$). This prevents us from extensive hyper-parameter tuning on each PDE. Please note that this lower the reported relative MSE, thus advantaging the baseline. Finally, for computational reasons, we did not fitted the PINNs for more than $1,500$ epochs depending on the dataset. This can explain the observed performances of this baseline since PINNs are traditionally trained for more than $10,000$ epochs. We detail in table 4 the hyper-parameters for each dataset.
Hyper-parameter Dataset
1d 1d + time 2d 2d + time
Helmholtz Poisson NLRD Darcy Heat
epochs $1,000$ $1,000$ $1,500$ $1,500$ $1,000$
learning rate 1e-4 1e-5 1e-4 1e-3 1e-3
Table 4: Hyper parameters for PINNs+L-BFGS baseline.
Moreover, we used a $3$ layer NN with $256$ neurons and a GeLU activation function.
Basis configuration:
For all datasets, we use Spline of degree $3$, built with shifted nodes. We change the number of terms in the basis depending on the problem considered. For the $1d$-problem, we use $32$ terms. For the $2d$-problem, $40$ elements are in the basis of each dimension, except for Reaction-Diffusion where the variable t has $20$ terms. Moreover, during training, we use the projection of the initial conditions and/or parameters and/or forcings terms function in the basis as input to the networks.
Appendix D Dataset details
@@ D.1 Helmholtz
The solution can be analytically derived: $u(x)=\alpha\cos(\omega x+\beta)$, with $\beta=\arctan(\frac{-v_{0}}{\omega u_{0}})$, $\alpha=\frac{u_{0}}{\cos(\beta)}$ and directly computed from the PDE data. We generate $1,024$ trajectories for training and $256$ for testing with $u_{0},v_{0}\sim\mathcal{N}(0,1)$ and $\omega\sim\mathcal{U}(0.5,50)$ and compute the solution on $[0,1]$ with a spatial resolution of $256$. For training, we keep $800$ samples and use the complete dataset for the additional experiments presented in figs. 12(a) and 12(b). Moreover, we sub-sample the spatial resolution by $4$ and keep $64$ points.
@@ D.2 Poisson
We chose $f$ to be a non-linear forcing terms: $f(x)=\frac{\pi}{K}\sum_{i=1}^{K}a_{i}i^{2r}\sin(\pi x)$, with $a_{i}\sim\mathcal{U}(-100,100)$, we used $K=16$, $r=-0.5$, and solve the equation using a backward finite difference scheme. We generate $1,000$ trajectories with $u_{0},v_{0}\sim\mathcal{N}(0,1)$ and compute the solution on $[0,1]$ with a spatial resolution of $64$.
@@
Reaction-Diffusion with initial conditions
: To complexify the setting, we also change initial condition of the problem (NLRDIC in the following). The initial condition express as follows:
@@
D.3 Darcy Flow
The $2d$ Darcy Flow dataset is taken from (Li et al., 2020) and commonly used in the operator learning literature (Li et al., 2023b; Goswami et al., 2022).
@@ D.3 Darcy Flow
D.4 Heat
As a proof that our method can handle $2$d + time, we consider the dataset proposed by (Zhou & Farimani, 2024).
@@ D.4 Heat
Where $L=2$, $\nu$ is randomly chosen between $[2\times 10^{-3},2\times 10^{-2}]$, $A_{j}$ in $[-0.5,0.5]$, $l_{xj},l_{xy}$ are integers in $\{1,2,3\}$ and $\phi$ is in $[0,2\pi]$. As a difference with (Zhou & Farimani, 2024), we randomly chose $J$ between $1$ and $J_{max}=5$ to have more diversity in the represented frequencies in the data. The PDE are sampled with spatial resolution $64$ in x and y and temporal resolution $100$. However, during training these spatial resolution and subsampled by $4$ and coordinates are re scaled between 0 and 1. As for other PDEs, we use $800$ trajectories for training and $200$ for testing.
@@
D.5 Additional dataset: Advection
@@
D.6 Summary of problem settings considered
A summary of the datasets and parameters changing between $2$ trajectories are presented in table 5.
@@ E.1 Ablation
We experimentally show some properties of our iterative method. These evaluations are made on a test set composed of several instances of PDE with varying configurations $(\gamma,f,g)$. These experiments are performed on the Helmholtz equation which appeared as one of the more complex to optimize in our evaluation with $L=5$ unless stated otherwise.
@@
On fig. 11, we show that having more optimizer steps allows for a better generalization. However, we observed in our experiments that the generalization error stabilizes or even increases after the proposed $5$ steps. This limitation of the solver should be investigated in future work in order to allow the model to make more iterations. Note that for this experiment, we lowered batch size (and adapted learning rate accordingly) when the number of steps increased (fig. 11).
@@
On figs. 12(a) and 12(b), we show that compared to other physics-informed baselines, our solver requires fewer data to learn to solve PDE. Note that on this simple example, the MLP+basis baseline also performed well. However, as shown in table 2, this is not the case for all the datasets. The contribution of the iterative procedure clearly appears with 2×2$\times$2 × less data, our models performs best than this baseline.
@@
Table 6 shows that optimizing our network with physical loss greatly complexifies training. Indeed, adding an ill-conditioned loss to the standard MSE makes training ill-conditioned.
@@
Table 6: How the physical loss complexify training and lowers performances. Metrics in Relative MSE on the test set of Helmholtz equation.
@@
In table 7, we compare two optimizer configurations. "Direct" means that the network directly predicts the parameters for the next step (i.e. $\theta^{k+1}=\mathcal{F}_{\varrho}(\nabla\mathcal{L}_{\textnormal{PDE}}(\theta^{k}),\;\gamma,f,g)$), while "GD", corresponds to the update rule described before and using SGD as the base algorithm $\theta^{k+1}=\theta^{k}-\eta\mathcal{F}_{\varrho}(\nabla\mathcal{L}_{\textnormal{PDE}}(\theta^{k}),\;\gamma,f,g)$. The latter clearly outperforms the direct approach and shows that leaning increments is more efficient than learning a direct mapping between two updates. As shown in table 7 increasing the number of steps improves the performance (shown here for 1 and 5 update steps). However the performance does not improve anymore after a few steps (not shown here).
@@
On table 8, we study the performance of our proposed method with different inner learning rates $\eta$. As expected, a higher learning rate leads to better performances since, the optimization is taking bigger steps.
@@
As indicated in eq. 9 the inputs of our learned solver are $(\gamma,f,g,\nabla\mathcal{L}_{\textnormal{PDE}}(\theta_{L}))$. We performed experiments by removing either $\gamma$ or $\nabla\mathcal{L}_{\textnormal{PDE}}(\theta_{l})$ from the input (For the Helmholtz equation, there is no forcing term $f$). The BC $g$ are kept since they are part of the PDE specification and are required for ensuring the uniqueness of the solution. This experiment (table 9) illustrates that conditioning on the PDE parameters $\gamma$ is indeed required to solve the parametric setting. Without $\gamma$, the solver has no hint on which instance should be used. The addition of the gradient information, $\nabla$, which is at the core of our method, is indeed crucial for improving the convergence and validates our setting.
@@
Non-linear Reconstruction Finally, we showed by a simple experiment that our model can handle non linear cases (see table 10). We propose to model the solution $u$ using a non-linear combination of the basis terms $\phi_{i}$. The relation between the $\phi_{i}$s is modeled using a simple NN with one hidden layer and a tanh activation function. This experiment is performed on the Poisson PDE.

### v3

@@ Discussion
Appendix C Dataset details
@@ C.1 Helmholtz
The solution can be analytically derived: $u(x)=\alpha\cos(\omega x+\beta)$, with $\beta=\arctan(\frac{-v_{0}}{\omega u_{0}})$, $\alpha=\frac{u_{0}}{\cos(\beta)}$ and directly computed from the PDE data. We generate $1,024$ trajectories for training and $256$ for testing with $u_{0},v_{0}\sim\mathcal{N}(0,1)$, and $\omega\sim\mathcal{U}(0.5,50)$ and compute the solution on $[0,1]$ with a spatial resolution of $256$. For training, we keep $800$ samples and use the complete dataset for the additional experiments presented in figs. 12(a) and 12(b). Moreover, we sub-sample the spatial resolution by $4$ and keep $64$ points for training.
@@ C.2 Poisson
We chose $f$ to be a non-linear forcing term: $f(x)=\frac{\pi}{K}\sum_{i=1}^{K}a_{i}i^{2r}\sin(\pi x)$, with $a_{i}\sim\mathcal{U}(-100,100)$, we used $K=16$, $r=-0.5$, and solve the equation using a backward finite difference scheme. We generate $1,000$ trajectories with $u_{0},v_{0}\sim\mathcal{N}(0,1)$ and compute the solution on $[0,1]$ with a spatial resolution of $64$.
@@
Reaction-Diffusion with initial conditions:
To complexify the setting, we also change the initial condition of the problem (NLRDIC in the following). The initial condition is expressed as follows:
@@
C.3 Darcy Flow
The $2d$ Darcy Flow dataset is taken from (Li et al., 2020) and is commonly used in the operator learning literature (Li et al., 2023b; Goswami et al., 2022).
@@ C.3 Darcy Flow
C.4 Heat
As proof that our method can handle $2$d + time, we consider the dataset proposed by (Zhou & Farimani, 2024).
@@ C.4 Heat
Where $L=2$, $\nu$ is randomly chosen between $[2\times 10^{-3},2\times 10^{-2}]$, $A_{j}$ in $[-0.5,0.5]$, $l_{xj},l_{xy}$ are integers in $\{1,2,3\}$ and $\phi$ is in $[0,2\pi]$. As a difference with (Zhou & Farimani, 2024), we randomly chose $J$ between $1$ and $J_{max}=5$ to have more diversity in the represented frequencies in the data. The PDEs are sampled with spatial resolution $64$ in x and y and temporal resolution $100$. However, during training the spatial resolution are subsampled by $4$ and the coordinates are re scaled between 0 and 1. As for other PDEs, we use $800$ trajectories for training and $200$ for testing.
@@
C.5 Additional dataset: Advection
@@
C.6 Summary of problem settings considered
A summary of the datasets and parameters changing between $2$ trajectories is presented in table 3.
@@ C.6 Summary of problem settings considered
Appendix D Implementation details
D.1 B-Spline basis
Figure 10: B-spline basis with $N=10$ terms with shifted spline (Left) and higher multiplicity nodes (Right). Dashed lines represent nodes’ position with color the darker, the higher the multiplicity.
Higer-dimension basis
For $1$d + time, $2$d dataset, and $2$d + time we build a $2$ (or $3$) dimensional B-spline basis, i.e., we treat the time coordinates as spatial ones. To build such bases, we compute the Cartesian product between the $2$ (or $3$) $1$d-bases, $1$ per dimension. This means that for a $2$d dataset, for which we chose to use bases of size $N1$ and $N2$ for each coordinate, the resulting basis will have $N1+N2+N1*N2$ terms. This method makes the training more costly and several techniques to improve its scalability could be used. For $3$d datasets, the number of terms in the basis is cubic.
D.2 Training details
In our experiments, neural networks are trained using the Adam optimizer. For network optimization, we employ a smooth $l1$-loss for our solver while for other baselines, we use MSE loss and/or physical losses. All models are trained for at least $1,500$ epochs on datasets composed of some sampling of $\gamma$ and/or $f$ and/or $g$. If not stated otherwise, we train our proposed method for $750$ epochs and baselines for $1,500$ epochs. We use the Adam optimizer and an initial learning rate of $0.001$. We use an exponential learning rate scheduler that lowers the learning rate by $0.995$ every epoch. Experiments are conducted on NVIDIA TITAN V ($12$ Go) for $1$d datasets and NVIDIA RTX A6000 GPU with 49Go for $1d$ + time, $2$d datasets or $2$d + time. We recall in algorithm 3 the pseudo-code for training our proposed method.
Data: $\Theta_{0}\in\mathbb{R}^{n}$, PDE ($\gamma,f,g)$, sample values $u(x)$
Result: $\mathcal{F}_{\varrho}$
for $e=1...$ epochs do
       for (PDE, x, u) in dataset do
            Initialize $\Theta_{0}$
             Estimate $\Theta_{L}$ from $\Theta_{0},(\gamma,f,g)$ using Algorithm 1
             Reconstruct $u_{\Theta_{L}}(x)$
       end for
end for
return $\mathcal{F}_{\varrho}$
D.3 Models
We present here the training details for our model and the baseline. The results can be found in table 2. In all our experiments, we use a GeLU (Hendrycks & Gimpel, 2016) activation function. The details of the model architecture on each dataset are presented in table 4.
Model Architecture Dataset
1d 1d + time 2d 2d + time
Helmholtz Poisson NLRD Darcy Heat
PINNs MLP depth 3 3 3 3 3
MLP width 256 256 256 256 256
PPINNs MLP depth 8 3 5 3 3
MLP width 64 256 256 256 256
PO-DeepONet Branch Net depth 3 2 5 5 5
Branch Net width 256 256 256 256 256
Trunk Net depth 3 2 5 5 5
Trunk Net width 256 256 256 256 256
P2INNs Enc params depth 4 4 4 4 4
Enc params width 256 256 256 256 256
Emb params 128 128 128 128 128
Enc coord depth 3 3 3 3 3
Enc coord width 256 256 256 256 256
Emb coord 128 128 128 128 128
Dec depth 6 6 6 6 6
Dec width 256 256 256 256 256
Activations GeLU GeLU GeLU GeLU GeLU
MLP + basis MLP depth 5 5 5 5 5
MLP width 256 256 1,024 1,024 1,024
PI-DeepONet Branch Net depth 3 2 5 5 5
Branch Net width 256 256 256 256 5
Trunk Net depth 3 2 5 5 5
Trunk Net width 256 256 256 256 256
PINO FNO depth 3 3 4 3 3
FNO width 64 64 64 64 64
FNO modes 1 16 16 10 20 7
FNO modes 2 - - 5 20 7
FNO modes 3 - - - - 5
FNO fc dim 64 64 64 64 64
Ours FNO depth 3 3 3 3 3
FNO width 64 64 64 64 64
FNO modes 1 16 16 10 20 7
FNO modes 2 - - 5 20 7
FNO modes 3 - - - - 5
FNO fc dim 64 64 64 64 64
Table 4: Architecture details of our model and baselines
Hyperparameters for Helmholtz:
Training of PPINNs:
For the Helmholtz dataset, we trained our model for 2000 epochs with a plateau learning rate scheduler with patience of 400 epochs. For the Poisson dataset, we trained our parametric PINN model for 5000 epochs with a cosine annealing scheduler with a maximum number of iterations of 1000. For the Reaction-diffusion dataset, we consider an initial learning rate of $0.0001$ instead of $0.001$. Finally, P2INNs is trained for $5,000$ epochs using the Adam optimizer and an exponential scheduler with patience $50$.
PINNs baselines (PINNs + L-BFGS and PINNs-multi-opt):
The conditioning of the problem highly depends on the parameters of the PDE and initial/boundary conditions. This can lead to unstable training when optimizing PINNs, requiring specific parameter configurations for each PDE. To avoid extensive research of the best training strategy, we found a configuration that allowed a good fitting of most of the testing dataset. This means that in the values reported in table 2, we removed trainings for which the losses exploded (only a few hard PDEs were removed, typically between none to $20$). This prevents us from extensive hyper-parameter tuning on each PDE. Please note that this lowers the reported relative MSE, thus advantaging the baseline. We detail in tables 5 and 6 the hyper-parameters for each dataset.
Hyper-parameter Dataset
1d 1d + time 2d 2d + time
Helmholtz Poisson NLRD Darcy Heat
epochs $1,000$ $1,000$ $1,500$ $1,500$ $1,000$
learning rate 1e-4 1e-5 1e-4 1e-3 1e-3
Table 5: Hyperparameters for PINNs+L-BFGS baseline.
Hyper-parameter Dataset
1d 1d + time 2d 2d + time
Helmholtz Poisson NLRD Darcy Heat
epochs Adams $800$ $800$ $800$ $1,200$ $1,200$
epochs L-BFGS $200$ $200$ $200$ $300$ $300$
epochs total $1,000$ $1,000$ $1,000$ $1,500$ $1,500$
learning rate Adam 1e-4 1e-5 1e-4 1e-3 1e-3
learning rate L-BFGS 1 1 1e-3 1 1
Table 6: Hyperparameters for PINNs-multi-opt baseline.
Basis configuration:
For all datasets, we use Splines of degree $3$, built with shifted nodes. We change the number of terms in the basis depending on the problem considered. For the $1d$-problem, we use $32$ terms. For the $2d$-problem, $40$ elements are in the basis of each dimension, except for Reaction-Diffusion where the variable t has $20$ terms. Moreover, during training, we use the projection of the initial conditions and/or parameters and/or forcing terms function in the basis as input to the networks. Finally, for experiments using the Heat dataset (i.e. 3d basis), we used $15$ terms for the $x$ and $y$ spatial coordinates and $10$ terms for the $t$-coordinate.
@@ E.1 Ablation
We experimentally show some properties of our iterative method. These evaluations are made on a test set composed of several instances of PDE with varying configurations $(\gamma,f,g)$ that are unseen during training. These experiments are performed on the Helmholtz equation which appeared as one of the most complex to optimize in our evaluation. We used $L=5$ in our method for a better visualization, unless stated otherwise.
@@
On fig. 11, we show that having more optimizer steps allows for a better generalization. However, we observed in our experiments that the generalization error stabilizes or even increases after the proposed $5$ steps. This limitation of the solver should be investigated in future work in order to allow the model to make more iterations. Note that for this experiment, we lowered batch size (and adapted the learning rate accordingly) when the number of steps increased (fig. 11).
@@
On figs. 12(a) and 12(b), we show that compared to other physics-informed baselines, our solver requires less data to learn to solve PDE. Note that in this simple example, the MLP+basis baseline also performed well. However, as shown in table 2, this is not the case for all the datasets. The contribution of the iterative procedure clearly appears since with 2×2$\times$2 × less data, our model performs better than this baseline.
@@
Table 7 shows that optimizing our network with physical loss greatly complicates training. Indeed, adding an ill-conditioned loss to the standard MSE makes training ill-conditioned.
@@
Table 7: How the physical loss complexifies training and lowers performances. Metrics in Relative MSE on the test set of Helmholtz equation.
@@
In table 8, we compare two optimizer configurations. "Direct" means that the network directly predicts the parameters for the next step (i.e. $\Theta_{k+1}=\mathcal{F}_{\varrho}(\nabla\mathcal{L}_{\textnormal{PDE}}(\Theta_{k}),\gamma,f,g)$), while "GD", corresponds to the update rule described before and using SGD as the base algorithm $\Theta_{k+1}=\Theta_{k}-\eta\mathcal{F}_{\varrho}(\nabla\mathcal{L}_{\textnormal{PDE}}(\Theta_{k}),\;\gamma,f,g)$. The latter clearly outperforms the direct approach and shows that learning increments is more efficient than learning a direct mapping between two updates. As shown in table 8 increasing the number of steps improves the performance (shown here for 1 and 5 update steps). However, the performance does not improve anymore after a few steps (not shown here).
@@
In table 9, we study the performance of our proposed method with different inner learning rates $\eta$. As expected, a higher learning rate leads to better performances since the optimization is taking bigger steps.
@@
As indicated in eq. 9 the inputs of our learned solver are $(\gamma,f,g,\nabla\mathcal{L}_{\textnormal{PDE}}(\theta_{L}))$. We performed experiments by removing either $\gamma$ or $\nabla\mathcal{L}_{\textnormal{PDE}}(\theta_{l})$ from the input (For the Helmholtz equation, there is no forcing term $f$). The BC $g$ are kept since they are part of the PDE specification and are required to ensure the uniqueness of the solution. This experiment (table 10) illustrates that conditioning on the PDE parameters $\gamma$ is indeed required to solve the parametric setting. Without $\gamma$, the solver has no hint on which instance should be used. The addition of the gradient information, $\nabla$, which is at the core of our method, is also crucial for improving the convergence and validates our setting.
@@
We showed by a simple experiment that our model can handle nonlinear cases (see table 11). We propose to model the solution $u$ using a non-linear combination of the basis terms $\phi_{i}$. The relation between the $\phi_{i}$ is modeled using a simple NN with one hidden layer and a tanh activation function. This experiment is performed on the Poisson PDE.
@@
Network architecture We show in table 12, an ablation on the layer type used for $\mathcal{F}_{\varrho}$ in our experiments: MLP, Residual Network (ResNet), FNO and a modified version of a MLP (ModMLP) taken from (Wang et al., 2021a), inspired from attention mechanism. We conducted this experiment on the Helmholtz dataset.
Table 12: Ablation on different layer types. Metric on the test set of the Helmholtz equation.
Layer type Relative MSE
MLP 8.25e-1
ResNet 6.87e-1
ModMLP 5.55e-2
FNO 2.41e-2
Irregular grids We show in table 13, an ablation on different types of grids: regular as in table 2, and irregular. The latter were created by sampling uniformly 75% of the points in the original grid. The sampled grids are different between each trajectory both during the training and testing phases. We conducted this experiment on the Helmholtz dataset. Finally, we show some reconstruction examples in figs. 13(a) and 13(b).
Table 13: Comparison of the performances when training our solver using regular or irregular and different grids for each PDE. Metrics on the test set.
Grid Relative MSE
regular 2.41e-2
irregular 3.38e-1
Figure 13: Comparison of the reconstruction of the solution when models are trained on regular (left) and irregular grid (right).
This experiment shows that our method is capable of handling irregular grids. We observe a decrease in the performances (table 13) and in the reconstruction quality figs. 13(a) and 13(b). However, we also note that our method is still capable of reconstructing the dynamic of the PDE, where most of the considered baselines failed to capture the oscillation in the Helmholtz PDE solution.
Error as a function of the PDE parameters values We illustrate in fig. 14 the behavior of the reconstruction of the MSE varying the PDE parameter values. We conducted this experiment on the Helmholtz PDE and varied $\omega$ from $-5$ to $55$ i.e. extrapolation of $10\%$ beyond the parameter distribution (and kept fixed boundary conditions).
Figure 14: MSE comparison with PDE parameters $\omega$ of the Helmholtz dataset and extrapolation outside of the training distribution.
We observe in fig. 14 that, as expected, solving Helmholtz PDE outside the training distribution fails for our model as well as for all baselines. The proposed method performs well inside the training distribution (as already observed in table 2) and behaves similarly to fully supervised methods on Out-Of-Distribution (OOD) examples. Finally, other baselines, involving physical only or hybrid training, only predict a mean solution.